# Application of the microscopic optical potential of chiral effective field theory in astrophysical neutron-capture reactions

### v1

@@ Abstract
The microscopic global nucleon–nucleus optical potential proposed by Whitehead, Lim, and Holt (WLH) is a state-of-the-art potential developed within the framework of many-body perturbation theory using realistic nuclear interactions from chiral effective field theory. Given its potentially greater predictive power for reactions involving exotic isotopes, we apply it to the calculations of astrophysical neutron-capture reactions for the first time, which are particularly important to the nucleosynthesis of elements heavier than iron. It is found that this potential provides a good description of experimental known neutron-capture cross sections and Maxwellian-averaged cross sections. For unstable neutron-rich nuclei, we comprehensively calculate the neutron-capture reaction rates for all nuclei with $26\leq Z\leq 84$, located between the valley of stability and the neutron drip line, using the backward-forward Monte Carlo method with the $f_{rms}$ deviation as the $\chi^{2}$ estimator. The results reveal a noticeable separation in the uncertainty of rates around an isospin asymmetry of 0.28 under the constraint $f_{rms}\leq 1.56$. This highlights the critical role of isospin dependence in optical potentials and suggests that future developments of the WLH potential may pay special attention to the isospin dependence.
@@ 1 Introduction
The elements heavier than iron in the Universe are mostly synthesized through neutron-capture reactions [1, 2]. The slow ($s$) [3, 4, 5, 6] and rapid ($r$) [7, 8, 9, 10] neutron-capture processes with their neutron-capture timescales slower or more rapid, respectively, than $\beta$ decay timescale of unstable nuclei are responsible for the vast majority of heavy element abundances. Between these two processes, other neutron-capture processes, such as the intermediate ($i$) [11, 12, 13] and neutron ($n$) [14, 15] processes, may also be indispensable for explaining certain specific features in elemental abundance patterns. However, a large number of neutron-capture rates involved in these nucleosynthesis processes are not known experimentally [16, 17, 18], and only theoretical predictions can fill the gaps. Theoretical neutron-capture rates are commonly treated within the framework of the Hauser-Feshbach statistical model [19, 20]. The optical potential, nuclear level density, and $\gamma$-ray strength function are crucial nuclear structure inputs to this model. Significant efforts have been devoted to improving the descriptions of the nuclear level densities, both from phenomenological [21, 22, 23, 24, 25] and microscopic [26, 27, 28] perspectives, and the strength functions through phenomenological [29, 30, 31, 32, 33] and microscopic [34, 35, 36] approaches.
@@
For the optical potential, which is essential for calculating the neutron transmission coefficient, a large number of developed models are available [37]. These include phenomenological potentials [38], semi-microscopic potentials [39, 40, 41], dispersive potentials [42], as well as models with uncertainty quantification [43, 44, 45]. However, most of them either rely on fitting nuclear reaction data, introducing unreliability when extrapolated to exotic nuclei lacking experimental constraints, or have limited applicability ranges, preventing their global use in astrophysical neutron-capture reactions [46]. The microscopic optical potentials developed from chiral effective field theory is a new booming area [47, 48, 49, 50, 51]. Recently, Whitehead, Lim, and Holt (WLH) constructed the full microscopic global nucleon-nucleus optical potential based on an analysis of 1800 isotopes within the framework of many-body perturbation theory using realistic nuclear interactions from chiral effective field theory [52]. This state-of-the-art potential, derived from the chiral nuclear force without any parameters fitted to nucleon-nucleus scattering data, is expected to be more powerful in predicting neutron-capture rates for nuclei far from stability. Moreover, it provides quantified uncertainties for the parameters, enabling one to estimate the uncertainty in a given reaction observable. While the WLH potential displays impressive capability in describing elastic scattering [52], charge-exchange reactions [53], and transfer reactions [54], it has not yet been widely applied to astrophysical neutron-capture reactions.
@@
In this work, we apply the WLH potential to the study of astrophysical neutron-capture reactions for the first time and test its applicability to nuclei with $26\leq Z\leq 84$ lying between the neutron drip line and the valley of $\beta$ stability, which is a critical region for the nucleosynthesis of elements heavier than iron. The rest of this article is organized as follows. In Sec. 2, the nuclear theoretical models used in our work are described. In Sec. 3, the calculated results are compared with experimental data, and the WLH potential is applied to neutron-rich nuclei. In Sec. 4, the conclusions are presented.
2 Theoretical model
We use the Hauser-Feshbach statistical model to calculate the neutron-capture cross sections and therefore the reaction rates. This model is based on the fundamental Bohr hypothesis that the capture process occurs by means of the intermediary production of a compound system that can reach a state of thermodynamic equilibrium. Provided the reaction $A+n\rightarrow B^{x}+\gamma$ represents the neutron capture onto the nucleus $A$ leaving the residual nucleus $B$ in the $x$th excited state and the photon, the corresponding cross section can be expressed as [20]
$\sigma_{A+n\rightarrow B^{x}+\gamma}(E)=$ $\frac{\pi}{k^{2}}\sum^{l_{max}+I_{A}+I_{n}}_{J=mod(I_{A}+I_{n},1)}\sum^{1}_{\Pi=-1}\frac{2J+1}{(2I_{A}+1)(2I_{n}+1)}$
×$\times\sum^{J+I_{A}}_{J_{n}=\left|J-I_{A}\right|}\sum^{J_{n}+I_{n}}_{l_{i}=\left|J_{n}-I_{n}\right|}\sum^{J+I^{x}_{B}}_{\lambda=\left|J-I^{x}_{B}\right|}\sum^{\lambda+I_{\gamma}}_{l_{f}=\left|\lambda-I_{\gamma}\right|}\delta^{\pi}_{C_{n}}\delta^{\pi}_{C_{\gamma}}$
×$\times\frac{\left\langle T^{J}_{C_{n},l_{i},J_{n}}(E)\right\rangle\left\langle T^{J}_{C_{\gamma},l_{f},\lambda}(E_{\gamma})\right\rangle}{\sum_{Clj}\delta^{\pi}_{C}\left\langle T^{J}_{C,l,j}(E_{C})\right\rangle}W^{J}_{C_{n}l_{i}J_{n}C_{\gamma}l_{f}\lambda}.$ (1)
In the above equations, $E$ ($E_{\gamma}$), $I_{n}$ ($I_{\gamma}$), $\pi_{n}$ ($\pi_{\gamma}$), $l_{i}$ ($l_{f}$), and $J_{n}$ ($\lambda$) represent the energy, spin, parity, orbital, and total angular momentum for the projectile neutron (emitted photon), respectively, and $l_{max}$ is the maximum value of $l_{i}$. The symbols $I_{A}$, $\pi_{A}$ ($I_{B}^{x}$, $\pi_{B}^{x}$) represent the spin and parity of the target (residual) nucleus, while $J$ and $\Pi$ correspond to the spin and parity of the compound system. The initial system ($A+n$) is designated by the channel label $C_{n}=(n,I_{n},E,E_{A}=0,I_{A},\pi_{A})$, and $C_{\gamma}=(\gamma,I_{\gamma},E_{\gamma},E_{B}^{x},I_{B}^{x},\pi_{B}^{x})$ is similar for the final system ($B^{x}+\gamma$), where $E_{A}$ and $E_{B}^{x}$ are the excited energy for target and residual nucleus, respectively. $\delta^{\pi}_{C_{n}}=1$ if $\pi_{A}\pi_{n}(-1)^{l_{i}}=\Pi$ and 0 otherwise. $\delta^{\pi}_{C_{\gamma}}=1$ if $\pi^{x}_{B}\pi_{\gamma}(-1)^{l_{f}}=\Pi$ and 0 otherwise. $k$ is the wave number of the relative motion. $W$ is the width fluctuation correction factor. $T$ is the transmission coefficient and ${\sum_{Clj}\delta^{\pi}_{C}\left\langle T^{J}_{C,l,j}(E_{C})\right\rangle}$ is the sum of transmission coefficient $T$ for all possible decay channels.
Each transmission coefficient $T$ is calculated for all levels with known energy, spin, and parity. If the excitation energy $E^{x}$, which is implicit in the definition of the channel $C$, corresponds to a state in the continuum, an effective transmission coefficient for an excitation-energy bin of width $\Delta E$ is defined by the integral
[…]
over the nuclear level density $\rho(E,J,\Pi)$. $T$ relies on the optical potentials of the two reacting particles. In this work, we use the WLH optical potential from Ref. [52] to calculate $T$ for the neutron. The parameters of the WLH potential are sampled within their respective uncertainty ranges, and the corresponding reaction observables are computed for each sample. The resulting uncertainties in the calculated observables are quantified through extensive sampling and computations. The calculations are performed with the reaction code TALYS [55].
@@ 3.1 Comparison with experimental result
Figure 1: Comparison of the calculated and experimental neutron-capture cross sections as a function of incident energy $E_{n}$ for three nuclei. The shaded area represents cross sections calculated from 400 random samples of the WLH potential. Experimental data are shown as black points and are taken from Refs. [56] for ${}^{61}\text{Ni}(\text{n},\gamma)^{62}\text{Ni}$, [57, 58] for ${}^{122}\text{Sn}(\text{n},\gamma)^{123}\text{Sn}$, and [59, 60, 61] for ${}^{176}\text{Lu}(\text{n},\gamma)^{177}\text{Lu}$.
@@
To further assess the global predictive ability of the WLH potential against all available experimental data, we calculate the Maxwellian-averaged cross sections (MACS) at 30 keV for 221 nuclei with $26\leq Z\leq 83$, and compare the results with experimental data taken from the KADoNiS database [62]. The average values of the WLH potential parameter set are used in the calculations. Other model parameters, such as those for level density and photon strength functions, are taken, wherever possible, from the recommended values in Ref. [63], which are suggested for TALYS calculations in astrophysics applications. As shown by the red circles in Fig. 2, the deviations between the MACS values calculated using the WLH potential and those reported in KADoNiS are generally within a factor of two. Furthermore, we present the theoretical results obtained using the phenomenological potential of Koning and Delaroche (KD) which is widely employed in nuclear reaction calculations and is well known for its good agreement with experimental data through parameter adjustment [38]. It can be seen in Fig. 2 that the results based on the WLH potential are consistent with those obtained from the KD potential in both magnitude and overall trend. These observations confirm the capability of the WLH potential to reproduce experimental data for stable nuclei and support its application to to unstable neutron-rich nuclei.
Sampling of the potential parameter set introduces an uncertainty range in the calculated results. To quantify the predictive accuracy of the WLH potential for each sample, we adopt the root-mean-square (rms) criterion, i.e. $f_{rms}$ deviation, defined in Ref. [63, 64] as
@@
where $N_{e}$ is the number of known MACS from KADoNiS database, and $\left\langle\sigma_{\mathrm{th},i}\right\rangle$ and $\left\langle\sigma_{\exp,i}\right\rangle$ are the theoretical and experimentally MACS, respectively. With Eq. (3), each sampling corresponds to an $f_{rms}$ deviation value, which is always greater than or equal to 1. The deviation from 1 indicates, on average, how far the prediction deviates from the experimental data. In Fig. 3, the distribution of $f_{rms}$ deviations with respect to the 221 experimental ($n,\gamma$) MACS at 30 keV from the KADoNiS database is displayed. As can be seen, while the $f_{rms}$ deviations reach up to 2.16, most are below 1.74. These results are consistent with those reported in Refs. [63, 64], which are based on uncertainty analyses of neutron-capture reactions relevant to astrophysics. The fact that the uncertainties remain within a reasonable range further confirms the validity of the WLH potential in describing neutron-capture experimental data.
Figure 3: Distribution of $f_{rms}$ deviations with respect to the 221 experimental ($n,\gamma$) MACS for 400 samples of the parameter set.
@@ 3.2 Application to neutron-rich nuclei
Applying the WLH potential to neutron-capture reactions of nuclei far from stability, for which no experimental data are available, leads to increasingly intensive computations. This is due to the large number of neutron-rich nuclei involved, each requiring extensive sampling of potential parameters across a wide uncertainty range, along with the corresponding reaction calculations. For the purposes of efficiently propagating parameter uncertainties to calculated observables, as well as reducing computational cost, we adopt the backward-forward Monte Carlo (BFMC) method [64, 65, 66, 67]. This method relies on the sampling of the model parameters and the use of a generalized $\chi^{2}$ estimator to quantify the likelihood of each simulation with respect to a given set of experimental constraints. It consists primarily of two Monte Carlo (MC) steps: in the backward MC step, suitable parameter samples are selected based on experimental constraints; in the forward MC step, the selected parameter samples are used to calculate unmeasured quantities.
@@
to select the sampled potential parameter sets for which the $f_{rms}$ deviation is less than or equal to the critical value $f_{rms}^{crit}$. This ensures that the corresponding calculated observables are in closer agreement with the experimental data than those with $f_{rms}>f_{rms}^{crit}$. In Fig. 2, the error bars of the WLH potential are derived from potential parameters constrained by $f_{rms}\leq 1.56$. It can be seen that the uncertainty range of the WLH potential effectively encompasses the theoretical results of the KD potential.
@@
In the forward MC step, the parameter sets selected in the backward MC step are used to calculate the neutron-capture rates of neutron-rich nuclei. As shown in Fig. 4, the neutron-capture rates and their uncertainty ranges as a function of temperature for the Sn isotopes ${}^{122}\text{Sn}$, ${}^{132}\text{Sn}$, and ${}^{150}\text{Sn}$ are calculated using three $f_{rms}^{crit}$ values. As $f_{rms}^{crit}$ decreases, the uncertainty range gradually narrows, indicating that the $f_{rms}$ deviation effectively controls the uncertainty propagation through the selection of potential parameters. For ${}^{122}\text{Sn}$ and ${}^{132}\text{Sn}$, the uncertainty ranges are within one order of magnitude. For ${}^{150}\text{Sn}$, however, the differences in the reaction rates reach up to five orders of magnitude. Fig. 5 shows the neutron-capture rates of Sn isotopes as a function of neutron number $N$ at a temperature of $T=1$ GK. It is more clearly seen that the uncertainty in the rates decreases with decreasing $f_{rms}^{crit}$, and when $N$ exceeds approximately 87, the uncertainty increases dramatically and exhibits an odd-even effect.
@@
Furthermore, using the parameter sets constrained by $f_{rms}\leq 1.56$, we comprehensively calculate the neutron-capture rates at $T=1$ GK for all nuclei with $26\leq Z\leq 84$, located between the valley of stability and the neutron drip line. The ratio of maximum to minimum rates is shown in the chart of the nuclides in Fig. 6. There is a noticeable separation around an isospin asymmetry, i.e., $\delta=(N-Z)/A$, of 0.28, below which the uncertainty range in the reaction rate remains within one order of magnitude, and above which it mostly exceeds three orders of magnitude. A significant effect of the isospin asymmetry on the astrophysical neutron-capture reactions, especially for nuclei off stability line, has been already suggested [68] and recently corroborated by experimental evidence [69]. The separation related to isospin asymmetry in Fig. 6 further confirms this effect. Although the WLH potential parameterizes a global optical potential for each chiral force that depends smoothly on the projectile energy as well as the target nucleus mass number and isospin asymmetry, the isospin dependence of this potential may not be adequately accounted for in nuclei far from the valley of stability. Therefore, future studies focusing on the isospin dependence of the WLH potential in nuclei far from stability are suggested.

### v2

@@ Abstract
A state-of-the-art microscopic global nucleon-nucleus optical potential has been developed by Whitehead, Lim, and Holt (WLH) within the framework of many-body perturbation theory, incorporating realistic nuclear interactions derived from chiral effective field theory. Given its potentially greater predictive power for reactions involving exotic isotopes, we apply it to the calculations of astrophysical neutron-capture reactions for the first time, which are particularly important to the nucleosynthesis of elements heavier than iron. It is found that this potential provides a good description of experimental known neutron-capture cross sections and Maxwellian-averaged cross sections. For unstable neutron-rich nuclei, we comprehensively calculate the neutron-capture reaction rates for all nuclei with $26\leq Z\leq 84$, located between the valley of stability and the neutron drip line, using the backward-forward Monte Carlo method with the $f_{rms}$ deviation as the $\chi^{2}$ estimator. The results reveal a noticeable separation in the uncertainty of rates around an isospin asymmetry of 0.28 under the constraint $f_{rms}\leq 1.56$. This highlights the critical role of isospin dependence in optical potentials and suggests that future developments of the WLH potential may pay special attention to the isospin dependence.
@@ 1 Introduction
The elements heavier than iron in the Universe are mostly synthesized through neutron-capture reactions burbidge1957 , wallerstein1997a . The slow ($s$) busso1999 , pignatari2010 , kappeler2011 , lugaro2023 and rapid ($r$) arnould2007 , thielemann2011 , kajino2019 , cowan2021 neutron-capture processes with their neutron-capture timescales slower or more rapid, respectively, than $\beta$ decay timescale of unstable nuclei are responsible for the vast majority of heavy element abundances. Between these two processes, other neutron-capture processes, such as the intermediate ($i$) hampel2016 , denissenkov2017 , choplin2021 and neutron ($n$) meyer2000 , pignatari2018 processes, may also be indispensable for explaining certain specific features in elemental abundance patterns. However, a large number of neutron-capture rates involved in these nucleosynthesis processes are not known experimentally goriely1998 , goriely2021 , choplin2023 , and only theoretical predictions can fill the gaps. Theoretical neutron-capture rates are commonly treated within the framework of the Hauser-Feshbach statistical model hauser1952 , goriely2008 . The optical potential, nuclear level density, and $\gamma$-ray strength function are crucial nuclear structure inputs to this model. Significant efforts have been devoted to improving the descriptions of the nuclear level densities, both from phenomenological gilbert1965 , dilg1973 , Ignatyuk1979 , wang2019 , wang2020 and microscopic demetriou2001 , hilaire2006 , goriely2008a perspectives, and the strength functions through phenomenological Brink1957 , axel1962 , kopecky1990 , goriely2019c , wang2023a and microscopic goriely2002 , goriely2004 , xu2021 approaches.
@@
For the optical potential, which is essential for calculating the neutron transmission coefficient, a large number of developed models are available capote2009a . These include phenomenological potentials koning2003 , semi-microscopic potentials bauge1998a , bauge2000a , bauge2001a , dispersive potentials morillon2007a , as well as models with uncertainty quantification beyer2025 , pruitt2024a , pruitt2023b . However, most of them either rely on fitting nuclear reaction data, introducing unreliability when extrapolated to exotic nuclei lacking experimental constraints, or have limited applicability ranges, preventing their global use in astrophysical neutron-capture reactions hebborn2023a . The microscopic optical potentials developed from chiral effective field theory is a new booming area durant2022c , durant2022b , bai2021 , durant2020 , durant2018 . Recently, Whitehead, Lim, and Holt (WLH) constructed the full microscopic global nucleon-nucleus optical potential based on an analysis of 1800 isotopes within the framework of many-body perturbation theory using realistic nuclear interactions from chiral effective field theory whitehead2021 . This state-of-the-art potential, derived from the chiral nuclear force without any parameters fitted to nucleon-nucleus scattering data, is expected to be more powerful in predicting neutron-capture rates for nuclei far from stability. Moreover, it provides quantified uncertainties for the parameters, enabling one to estimate the uncertainty in a given reaction observable. While the WLH potential displays impressive capability in describing elastic scattering whitehead2021 , charge-exchange reactions whitehead2022 , and transfer reactions xu2024 , it has not yet been widely applied to astrophysical neutron-capture reactions.
@@
In this work, we apply the WLH potential to the study of astrophysical neutron-capture reactions for the first time and test its applicability to nuclei with $26\leq Z\leq 84$ lying between the neutron drip line and the valley of $\beta$ stability, which is a critical region for the nucleosynthesis of elements heavier than iron. The rest of this article is organized as follows. In Sec. 2, the WLH potential model used in our work is described. In Sec. 3, the calculated results are compared with experimental data, and the WLH potential is applied to neutron-rich nuclei. In Sec. 4, the conclusions are presented.
2 WLH potential model
In the WLH potential, a global optical potential is parameterized for each of the five chiral interactions with a Woods-Saxon form whitehead2021 ,
$U(r,E)=$ $-\mathcal{U}_{V}f(r;r_{V},a_{V})-i\mathcal{U}_{W}f(r;r_{W},a_{W})$
$+i4a_{S}\mathcal{U}_{S}\frac{d}{dr}f(r;r_{S},a_{S})$
$+\mathcal{U}_{SO}\frac{1}{m^{2}_{\pi}}\frac{1}{r}\frac{d}{dr}f(r;r_{SO},a_{SO})\vec{\ell}\cdot\vec{\sigma}.$ (1)
It consists of a real volume term $U_{V}$, an imaginary volume term $U_{W}$, an imaginary surface term $U_{S}$, and a real spin–orbit term $U_{SO}$. The Woods-Saxon shape factor is given by
$f(r;r_{i},a_{i})=\frac{1}{1+e^{(r-A^{1/3}r_{i})/a_{i}}},$ (2)
where $A$ is the mass number, and $r_{i}$ and $a_{i}$ are the geometry parameters .The well depths ($\mathcal{U}_{V}$, $\mathcal{U}_{W}$, $\mathcal{U}_{S}$, and $\mathcal{U}_{SO}$) and the geometry parameters depend smoothly on $A$, the projectile energy, and the isospin asymmetry, with only a few parameters. The functional forms of these parameters can be found in Ref. whitehead2021 .
The dominant source of theoretical uncertainty in the WLH potential arises from the choice of chiral potential. Therefore, in Ref. whitehead2021 , the covariance matrix for all the global optical potential parameters derived from five chiral interactions, along with their mean values, are used to generate a multivariate normal distribution that includes the correlations between parameters. Random parameter sets for the global optical potential can then be produced by sampling from this distribution. This enables one to estimate the uncertainty in a given reaction observable through many samples of the global optical potential. As a result, the multivariate distribution, and the uncertainty obtained from many samples within it, are independent of any experimental nucleon-nucleus reaction data, and the theoretical uncertainty stems directly from the uncertainty in the fundamental chiral interaction.
In our calculations, we sample parameter sets from the multivariate distribution which is implemented in Python code of Ref. whitehead2021 . The uncertainty range used for sampling is consistent with the original uncertainty ranges of the WLH parameters. The corresponding reaction observables for each sample are computed within the framework of the Hauser-Feshbach statistical model using the reaction code TALYS koning2023 .
@@ 3.1 Comparison with experimental result
Figure 1: Comparison of the calculated and experimental neutron-capture cross sections as a function of incident energy $E_{n}$ for three nuclei. The shaded area represents cross sections calculated from 400 random samples of the WLH potential. Experimental data are shown as black points and are taken from Refs. tomyo2005 for $^{61}\text{Ni}(\text{n},\gamma)^{62}\text{Ni}$, tolstikov1968 , timokhov1988 for $^{122}\text{Sn}(\text{n},\gamma)^{123}\text{Sn}$, and beer1984 , macklin1967 , wisshak2006 for $^{176}\text{Lu}(\text{n},\gamma)^{177}\text{Lu}$.
@@
To further assess the global predictive ability of the WLH potential against all available experimental data, we calculate the Maxwellian-averaged cross sections (MACS) at 30 keV for 221 nuclei with $26\leq Z\leq 83$, and compare the results with experimental data taken from the KADoNiS database dillmann2006a . The average values of the WLH potential parameter sets are used in the calculations. Other model parameters, such as those for level density and photon strength functions, are taken, wherever possible, from the recommended values in Ref. rochman2025 , which are suggested for TALYS calculations in astrophysics applications. As shown by the red circles in Fig. 2, the deviations between the MACS values calculated using the WLH potentials and those reported in KADoNiS are generally within a factor of two. Furthermore, we present the theoretical results obtained using the phenomenological potential of Koning and Delaroche (KD) which is widely employed in nuclear reaction calculations and is well known for its good agreement with experimental data through parameter adjustment koning2003 . It can be seen in Fig. 2 that the results based on the WLH potentials are consistent with those obtained from the KD potentials in both magnitude and overall trend. The consistency of the WLH results with the experimental MACS data, as well as with the KD results, demonstrates that the WLH potentials are capable of reproducing experimental neutron-capture data, though not with greater accuracy than KD.
Sampling of the potential parameter set introduces an uncertainty range in the calculated results. To quantify the predictive accuracy of the WLH potential for each sample, we adopt the root-mean-square (rms) criterion, i.e. $f_{rms}$ deviation, defined in Ref. rochman2025 , martinet2024 as
@@
where $N_{e}$ is the number of known MACS from KADoNiS database, and $\left\langle\sigma_{\mathrm{th},i}\right\rangle$ and $\left\langle\sigma_{\exp,i}\right\rangle$ are the theoretical and experimentally MACS, respectively. With Eq. (3), each sampling corresponds to an $f_{rms}$ deviation value, which is always greater than or equal to 1. The deviation from 1 indicates, on average, how far the prediction deviates from the experimental data. In Fig. 3, the distribution of $f_{rms}$ deviations with respect to the 221 experimental ($n,\gamma$) MACS at 30 keV from the KADoNiS database is displayed. As can be seen, while the $f_{rms}$ deviations reach up to 2.16, most are below 1.74. Even though the $f_{rms}$ deviations are generally small, they are consistent with those reported in Refs. rochman2025 , martinet2024 , which are based on uncertainty analyses of neutron-capture reactions relevant to astrophysics. Smaller $f_{rms}$ values mean that the calculated MACS is closer to the experimental data and further confirm the capability of the WLH potentials to reproduce neutron-capture experimental data. In other words, the small $f_{rms}$ values also support the fact that the low-energy neutron-capture cross sections of stable nuclei are not highly sensitive to the optical potentials.
The sensitivity of the neutron strength function to optical potentials is investigated as well. Using the 400 sets of parameters, we calculate the neutron strength functions for the S-wave ($S_{0}$) of 207 nuclei and the P-wave ($S_{1}$) of 91 nuclei, and compare them with the experimental data taken from RIPL-3 capote2009b , as shown in Fig. 4. It can be seen that the WLH potentials are globally compatible with the neutron strength data, but they span a much larger variation in the predicted strengths, which could be considered for further constraining the WLH potentials.
Figure 3: Distribution of $f_{rms}$ deviations. Blue bars show the deviations with respect to the 221 experimental ($n,\gamma$) MACS for 400 sampled parameter sets. Red bars correspond to the cases with $f^{\prime}_{frms}\leq 2.1$ with respect to the 298 experimental neutron strength functions.
Figure 4: Comparison between the experimental $S$- and $P$-wave strength functions ($S_{0}$ and $S_{1}$) for 298 nuclei ($26\leq Z\leq 83$) and the theoretical predictions. The red circles show the values calculated using 400 sets of WLH potential parameters. The blue shaded areas correspond to the parameter sets constrained by $f_{rms}\leq 1.56$.
@@ 3.2 Application to neutron-rich nuclei
Applying the WLH potential to neutron-capture reactions of nuclei far from stability, for which no experimental data are available, leads to increasingly intensive computations. This is due to the large number of neutron-rich nuclei involved, each requiring extensive sampling of potential parameters across a wide uncertainty range, along with the corresponding reaction calculations. For the purpose of reducing computational cost while efficiently propagating parameter uncertainties to the observables, as well as to investigate the properties of the WLH potential under different parameter constraint conditions, we adopt the backward-forward Monte Carlo (BFMC) method martinet2024 , chadwick2007 , bauge2011 , goriely2014 . This method relies on the sampling of the model parameters and the use of a generalized $\chi^{2}$ estimator to quantify the likelihood of each simulation with respect to a given set of experimental constraints. It consists primarily of two Monte Carlo (MC) steps: in the backward MC step, suitable parameter samples are selected based on experimental constraints; in the forward MC step, the selected parameter samples are used to calculate unmeasured quantities.
@@
to select the sampled potential parameter sets for which the $f_{rms}$ deviation is less than or equal to the critical value $f_{rms}^{crit}$. This ensures that the corresponding calculated observables are in closer agreement with the experimental data than those with $f_{rms}>f_{rms}^{crit}$. In Fig. 2, the error bars of the WLH potentials are derived from potential parameters constrained by $f_{rms}\leq 1.56$. It can be seen that the uncertainty range of the WLH potentials effectively encompasses the theoretical results of the KD potentials.
@@
In the forward MC step, the parameter sets selected in the backward MC step are used to calculate the neutron-capture rates of neutron-rich nuclei. As shown in Fig. 5, the neutron-capture rates and their uncertainty ranges as a function of temperature for the Sn isotopes $^{122}\text{Sn}$, $^{132}\text{Sn}$, and $^{150}\text{Sn}$ are calculated using three $f_{rms}^{crit}$ values. As $f_{rms}^{crit}$ decreases from 1.74 to 1.50, the uncertainty ranges clearly decrease, indicating that it is valid to use the experimental MACS to constrain the selection of parameter sets in the backward MC step, and that parameter sets closer to the experimental data are preferentially selected. For $^{122}\text{Sn}$ and $^{132}\text{Sn}$, the uncertainty ranges are within one order of magnitude. For $^{150}\text{Sn}$, however, the differences in the reaction rates reach up to five orders of magnitude. Fig. 6 shows the neutron-capture rates of Sn isotopes as a function of neutron number $N$ at a temperature of $T=1$ GK. It is more clearly seen that the uncertainty in the rates decreases with decreasing $f_{rms}^{crit}$, and when $N$ exceeds approximately 87, the uncertainty increases dramatically and exhibits an odd-even effect.
@@
Furthermore, using the parameter sets constrained by $f_{rms}\leq 1.56$, we comprehensively calculate the neutron-capture rates at $T=1$ GK for all nuclei with $26\leq Z\leq 84$, located between the valley of stability and the neutron drip line. The ratio of maximum to minimum rates is shown in the chart of the nuclides in Fig. 7. There is a noticeable separation around an isospin asymmetry, i.e., $\delta=(N-Z)/A$, of 0.28, below which the uncertainty range in the reaction rate remains within one order of magnitude, and above which it mostly exceeds three orders of magnitude. In addition, Fig. 6 for the Sn isotopes shows that the separation of uncertainties related to isospin asymmetry always occurs under the constraints of $f_{rms}^{crit}$ from 1.50 to 1.74, especially at 1.74, where almost no constraint is imposed. This implies that these separations are a general feature of the WLH potential and do not result from the constraint conditions. A significant effect of the isospin asymmetry on the astrophysical neutron-capture reactions, especially for nuclei off the stability line, has already been suggested goriely2007a and recently corroborated by experimental evidence voinov2021 . As pointed out in Ref. goriely2007a , the isovector imaginary component of the optical potential has a drastic impact on the neutron capture cross section for neutron-rich nuclei. Similarly, we believe that the separations observed in Figs. 6 and 7 may also be driven by the imaginary component associated with the isospin asymmetry in the WLH potential. For instance, from the expression of the imaginary volume term $U_{W}$ given in Ref. whitehead2021 , it can be seen that with increasing isospin asymmetry, $U_{W}$ may even take on negative values. This could be one of the factors responsible for the sharp increase in the uncertainties.
@@
Moreover, the calculation results for the strength functions obtained with the parameter sets constrained by $f_{rms}\leq 1.56$ are shown in Fig. 4, which exhibit improved predictions with reduced uncertainties. We also tentatively calculate the $f^{\prime}_{rms}$ deviations with respect to the 298 experimental neutron strength functions and find that nearly half of the parameter sets have $f^{\prime}_{rms}$ values below 2.1. The $f_{rms}$ distributions corresponding to the parameter sets with $f^{\prime}_{rms}\leq 2.1$ are displayed in Fig. 3. It can be seen that the $f_{rms}$ values associated with $f^{\prime}_{rms}\leq 2.1$ predominantly lie within the range $f_{rms}\leq 1.56$. These consistent constraints from the MACS and the strength functions indicate that similar conclusions to those presented in Figs. 57 would be reached when using either constraint. Ideally, employing the BFMC method with simultaneous constraints from both observables, MACS and strength functions, would be worthwhile in future work.